system of linear wquations. solve in substitution or addition method there is no book if possible express in ordered pair
3x-5y=1
10y=6x-1
thanks,
dnn

Answer by checkley71(8403)

3X-5Y=1
3X=5Y+1
X=5Y/3+1/3
THEN
10Y=6(5Y/3)+1/3
10Y=(30Y/3)+1/3
10Y=10Y+1/3
10Y-10Y=1/3
0=1/3
NO VALID SOLUTIONS FOR X OR Y